OpenStudy (anonymous):

Solve by addition or substitution. 4x-3y=24 8x-4y=56 Can someone tell me step by step how to do this I need to know for my test 2marrow.

OpenStudy (anonymous):

solve for x in 1st equation. x=(24+3y)/4, then input x on 2nd eq. 8(24+3y)/4-4y=56. then solve for Y and find its value. then input Y on fisrt equation and find X. hope it helps

OpenStudy (anonymous):

\[4x-3y=24[1]\] \[8x-4y=56[2]\] \[[1]\times 2\] \[2(4x-3y)=2(24)\] \[8x-6y=48[3]\] \[[2]-[3]\] \[(8x-4y)-(8x-6y)=(56)-(48)\] \[-4y--6y=8\] \[-4y+6y=8\] \[2y=8\] \[y=4\] To find x, sub in the y-value into one the three equations given by me above. (ie. [1], [2] or [3])

OpenStudy (anonymous):

what did you do when you got to (8x-4y)-(8x-6y)=(56)-(48) dont understand what u did to get ur answers

OpenStudy (anonymous):

I told you in that step I subtracted equation 3 from equation 2.

OpenStudy (anonymous):

[3]-[2]

OpenStudy (anonymous):

ok got it u canceled out 8x

OpenStudy (anonymous):

[2]-[3] sorry.

OpenStudy (anonymous):

Yeah. You have to get rid of a variable.

OpenStudy (anonymous):

That's the main thing about these. Getting rid of variables until you have one left.

OpenStudy (anonymous):

yeah i just didnt get it but now i see u canceled out 8x and it=-4y- -6

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Remember to find x.

OpenStudy (anonymous):

could you help me find x cause im confused

OpenStudy (anonymous):

I told you at the end.

OpenStudy (anonymous):

That's the most easiest part when you have one variable and you're finding the other.

OpenStudy (anonymous):

yeah that means x cud be 3 different numbers.

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

Try testing it for all three equations. You get the same x-value.

OpenStudy (anonymous):

no matter what happens, you get the same x-value...

OpenStudy (anonymous):

so how would I multiply 4 by 24?

OpenStudy (anonymous):

by using a calculator or getting better at multiplication. If you want to get better at multiplication I suggest you go back to grade 1.

OpenStudy (anonymous):

I ment is that what u want me to do for each one multiply it by the 24,56 and 48?

OpenStudy (anonymous):

no Where'd you get that from. The numbers in the brackets are just to tell you the number of equations I used. For example, [1] means the first equation, [2] means the second equation and [3] is the 3rd equation. What don't you get?

OpenStudy (anonymous):

Ok x=9 thanks.

OpenStudy (anonymous):

You're welcome.